OpenStudy (anonymous):

f(x) = x² g(x) = x +7 f(x)/ g(x) = ?

OpenStudy (aakashsudhakar):

f(x) = x^2 g(x) = x + 7 f(x)/g(x) = (x^2)/(x + 7) There's not really much you can do with this one.

OpenStudy (anonymous):

It's suppose to be simplified by subtracting exponents and reducing the variable. Which I think would make it x^2/7 but I'm not sure.

OpenStudy (anonymous):

I mean x^1 **